A series RLC circuit is connected to an oscillator with an rms voltage of 22.0 V, and consists of a 23.0 mH inductor, a 1.20 nF capacitor, and a 270. Ω resistor. If ω = ω0ω0, what is the power supplied to the circuit?
Answer #1

Since W=Wo (at resonance frequency ) current in the circuit

I=Vrms/R

=22.0/270

I=0.0815 A

So power supplied to the circuit

P=I2R

P=(0.0815)2X270

P=1.7934075

P=1.79W.